OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)) HELP? MEDAL AWARD

OpenStudy (anonymous):

f(x) = 9x - 2 g(x) = -x + 3 f(g(x)) = 9(g(x))-2 = 9(-x+3)-2 =-9x+27-2 =-9x+25

OpenStudy (anonymous):

i have a few more ?'s can u help me w them too

OpenStudy (anonymous):

let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)) ??

OpenStudy (anonymous):

i have to go fairly soon, ill be back on in about 3 hours if you are still on then. i can try to solve maybe one or two in the next 5 minutes

OpenStudy (anonymous):

that's all the time i need

OpenStudy (anonymous):

f(x) = x - 2 g(x) = x2 - 7x - 9 g(-1) = 2(-1) - 7(-1) - 9 g(-1) = -2 +7-9 g(-1) = -4 f(g(-1) = f(-4) f(-4) = -4 - 2 = -6

OpenStudy (anonymous):

so f(g(-1)) = -6

OpenStudy (anonymous):

-6 isnt an option -21 -3 3 21

OpenStudy (radar):

f(x) = x - 2 g(x) = x2 - 7x - 9 g(-1) = 2(-1) - 7(-1) - 9 Error is on this line. (-1)^2 =1 not -2 g(-1) = -2 +7-9 g(-1) = -4 f(g(-1) = f(-4) f(-4) = -4 - 2 = -6

OpenStudy (anonymous):

hmm, f(g(x)) = (g(x)) - 2 = 2x - 7x - 9 - 2 = -5x -11 = -5(-1) - 11 = 5-11 = -6 im not sure how this is so, could the book possibly be wrong?

OpenStudy (anonymous):

no lol

OpenStudy (anonymous):

oh, he meant to the power of 2, i didn't realize. i have to go know do u think u could help him from here for me, radar?

OpenStudy (anonymous):

i am a her. lol

OpenStudy (radar):

f(x) = x - 2 g(x) = x2 - 7x - 9 g(-1) = -1^2 - 7(-1) - 9 g(-1) =1 +7-9 g(-1) = -1

OpenStudy (radar):

f(x) = x - 2 f(g(-1) = -1 - 2 = -3

OpenStudy (radar):

@alwaysright13, did you see the error which gave a -6 that should be a -1??

OpenStudy (anonymous):

yes i did. what about this one ????Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (radar):

Just repeat the steps using the new function and the new value of 2. g(2)=2^2 - 6(2) -7. Do you see this is the g(x) functrion with 2 substituted for x?

OpenStudy (radar):

This now becomes g(2) = 4 - 12 - 7 = -15 or g(2) = -15 we now use this in the f(x) f(x)=x + 8 f(-15)=-15+8 = -7. Please post any additional problems as a new problem.